OpenStudy (anonymous):

Write the product in standard form (1+3i)(5-i)

OpenStudy (anonymous):

you have a choice you can use \((a+bi)(c+di)=(ac-bd)+(ad+bc)i\) or you can just multiply out like you would normally, remembering at the end that \(i^2=-1\)